OpenStudy (anonymous):

The area of a triangle is 24 square inches. What is the height of the triangle if the base length is 8 inches? 6 inches 8 inches 12 inches 16 inches

OpenStudy (campbell_st):

so \[24 = \frac{1}{2} \times 8 \times height\] solve for the height

OpenStudy (anonymous):

sooooo.......0.5*8 which equals 4.... so uhhhh

OpenStudy (campbell_st):

so \[24 = \frac{1}{2} \times 8 \times height\] solve for the height

OpenStudy (anonymous):

campbell is correct in her explanation. The answer is 6 inches.

OpenStudy (anonymous):

6 inches

OpenStudy (anonymous):

confusion O.0

OpenStudy (campbell_st):

so using your 4 times tables skills 4x ? = 24

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

I GET IT YAYAYAYAY

OpenStudy (campbell_st):

lol... glad to help..?

OpenStudy (anonymous):

THANK YOU DEARLY